# Connect Later: Improving Fine-tuning for Robustness with Targeted Augmentations

### v1

@@ 4 Connect Later: Pretrain First, Targeted Augmentations Later
Intuitively, pretraining learns good representations of the source and target domains and allows us to reuse the pretrained model for multiple downstream tasks, while targeted augmentations better connect the domains for the particular task.
@@ 4.1 Real-world examples of targeted augmentations
We design targeted augmentations for 4 real-world tasks: species identification from wildlife camera trap images, tumor detection, and astronomical time-series classification and redshift prediction. We show examples from the source, augmented, and target datasets for these tasks in Figure 2.
Figure 2: An example from the source dataset (left), an augmented version of the source example (middle), and an example from the target dataset (right) for our 3 tasks. (Top row) The target dataset in AstroClassification and Redshifts is much higher redshift than the source dataset. We apply the redshifting augmentation to simulate placing source objects at a higher redshift to better match the target dataset. The flux errors and flux values of the augmented example (middle) show much better resemblance of the target example. (Bottom row) The iWildCam-WILDS target dataset is composed of images from cameras unseen during fine-tuning, which are in potentially new habitats, so we randomize the habitat background by applying the Copy-Paste Same Y augmentation. This algorithm places source dataset animals into empty backgrounds from other cameras that have observed the same species. iWildCam-WILDS image examples shown here are from Gao et al. [17].
@@
For iWildCam-WILDS [18, 9], the task is to identify the wildlife species from static camera trap images. These cameras are placed in a wide variety of environments, which all have unique habitat conditions (e.g., African savannah vs. tropical rainforest) and camera characteristics (e.g., angles, resolutions). In this dataset, we use labeled data from 243 camera traps to learn a model that can generalize to data from 48 unseen camera traps.
@@
  • Targeted Augmentation: We augment the labeled dataset with the Copy-Paste Same Y algorithm, which uses image segmentation to copy-paste the animal onto different background images from cameras that have observed the same species [17].
@@
The task in Camelyon17-WILDS [19] is to classify whether a patch of a histopathology slide contains a tumor. These slides are contributed from multiple hospitals, which use different stain colors and also vary in distributions of patient cancer stage.
@@
  • Targeted Augmentation: We augment the labeled dataset with the Stain Color Jitter algorithm, which jitters the color of the slide image in the hematoxylin and eosin staining color space [29].
  • Task: Binary classification of whether a slide contains a tumor.
@@
Astronomical object classification [21, 30] involves predicting the object type (e.g., type II supernova) from a time series of an object’s brightness at multiple wavelengths (light curves). We curate this dataset from the Photometric LSST Astronomical Time Series Classification Challenge [PLAsTiCC, 20] (details in Appendix A.1).
@@
Similar to object type, redshift information is also available only for bright, nearby objects. We predict the scalar redshift value of each object and minimize mean squared error. This task has been studied for individual object types, such as quasars [31] and type Ia supernovae [32], but we consider a more realistic set of multiple object types. The labeled and unlabeled data are derived from the PLAsTiCC dataset. Redshifts is a new dataset that we contribute as part of this work.
@@ 4.2 Designing targeted augmentations
How do we design these targeted augmentations? We provide a general methodology based on matching the target distribution on a feature space:
@@
    Identify a feature space $\mathcal{Z}$. We assume that we can label $z\in\mathcal{Z}$ for each input and that the source and target domains largely differ on this feature space. One such example is the space of spurious, domain-dependent features (e.g., camera angle or resolution for iWildCam-WILDS), which is the approach followed by Gao et al. [17].
@@
We follow the procedure outlined above to design a targeted augmentation for AstroClassification and Redshifts, two astronomical time-series datasets (see Appendix B.2 for details). In these datasets, expert labels are only available for time-series of bright, nearby astronomical objects, while the unlabeled dataset contains mostly faint, distant objects. We describe distances in terms of cosmological redshift: nearby objects have lower redshift values than distant objects, causing the source and target redshift distributions to be mismatched (Appendix Figure 4).
@@
    The source and target domains primarily differ on their redshift distributions, so we identify this scalar feature as $z$.
@@
    We roughly fit the target redshift distribution while constraining the transformed redshift value to not be too far from the original redshift $z$, such that $\hat{p_{T}}(z^{\prime}\mid z)$ is distributed as $\text{loguniform}(0.95z,\;\text{min}(1.5(1+z)-1,\;5z))$, following Boone [21].
@@ 5 Experiments
We empirically test Connect Later with contrastive pretraining (iWildCam-WILDS, Camelyon17-WILDS) as well as pretraining with masked autoencoding (AstroClassification, Redshifts) to demonstrate Connect Later as a general fine-tuning method.
@@
For iWildCam-WILDS, we use a ResNet-50 model pretrained on unlabeled ImageNet data with SwAV contrastive learning [6]. We use a DenseNet121 pretrained on the unlabeled data provided in [9] with SwAV for Camelyon17-WILDS. To test Connect Later with other pretraining strategies, we use the masked autoencoding objective for AstroClassification and Redshifts by masking 60% of observations from each light curve (Appendix C). The same pretrained model is used for both tasks to demonstrate the reusability of pretrained features. We note that masked autoencoding has been linked to contrastive learning [33]. In particular, Zhang et al. [33] show that the masked autoencoding objective upper bounds the contrastive loss between positive pairs — thus, masked autoencoding implicitly aligns the positive pairs induced by the masking augmentations.
We fine-tune the pretrained models with linear probing then fine-tuning (LP-FT) [28], which has been shown to improve OOD performance.
@@
We evaluate our framework against three baselines: ERM, ERM+targeted augs, and standard fine-tuning. We include Avocado [21], the previous state-of-the-art model for AstroClassification. We also include a self-training baseline for AstroClassification and Redshifts, which has been shown to perform well on some real-world datasets [9]. For the self-training baseline, we pseudo-label the target dataset with a trained ERM+targeted augs model, then perform the same targeted augmentation on the pseudo-labeled target dataset. We then train a model with the pseudo-labeled and augmented target dataset combined with the labeled source dataset. We include additional domain adaptation baselines for iWildCam-WILDS and Camelyon17-WILDS: domain-adversarial neural networks (DANN) [11], correlation alignment (CORAL) [15], Noisy Student [34], and ICON [^2] https://github.com/a-tea-guy/ICON.
@@ 5.1 Main results
DANN  [9] $48.5\pm 3.2$ $31.9\pm 1.6$ $86.1\pm 1.3$ $64.5\pm 1.2$
CORAL  [9] $40.5\pm 1.6$ $27.9\pm 0.5$ $92.3\pm 0.7$ $62.3\pm 1.9$
Noisy Student  [9] $47.5\pm 1.0$ $32.1\pm 0.8$
@@
Tables 3 and 4 compare the results of Connect Later with baseline methods. On iWildCam-WILDS, Connect Later produces improvements in both ID and OOD performance despite the failure of standard fine-tuning. Connect Later substantially outperforms all other variants on the OOD metric, including state-of-the-art performances on AstroClassification by 3% OOD, iWildCam-WILDS by 0.8% OOD for ResNet-50, and Camelyon17-WILDS by 1.1% OOD for DenseNet121.
iWildCam.
@@
On iWildCam-WILDS, standard fine-tuning does not improve over ERM in ID performance and minimally improves over ERM in OOD performance, while ERM+targeted augmentations improves by 6% ID and OOD over both ERM and standard fine-tuning. While 3 of the domain adaptation baselines made ID and OOD improvements over standard fine-tuning (DANN: $46.4\%\rightarrow 48.5\%$ ID, $31.2\%\rightarrow 31.9\%$ OOD; Noisy Student: $46.4\%\rightarrow 47.5\%$ ID, $30.4\%\rightarrow 32.1\%$ OOD; ICON: $46.4\%\rightarrow 50.6\%$ ID, $30.4\%\rightarrow 34.5\%$), all fall short of ERM+targeted augs. Connect Later improves over both standard fine-tuning ($30.4\%\rightarrow 37.2\%$) and ERM+targeted augs ($36.3\%\rightarrow 37.2\%$) in OOD performance, achieving a new state-of-the-art performance for ResNet-50 on the iWildCam-WILDS benchmark.
Camelyon17.
@@
On Camelyon17-WILDS, standard fine-tuning produces significant gains over ERM in both ID ($89.3\%\rightarrow 92.3\%$) and OOD ($65.2\%\rightarrow 91.4\%$) average accuracy. ERM+targeted augmentations outperforms standard fine-tuning in ID accuracy ($92.3\%\rightarrow 96.7\%$), but does not improve OOD. DANN underperforms ERM in both ID and OOD accuracy, while CORAL produces similar ID accuracy as standard fine-tuning but poor OOD performance. Connect Later sets a new state-of-the-art on Camelyon17-WILDS with DenseNet121, improving on the best ID performance by 1.8% (ERM+targeted augs, $96.7\%\rightarrow 98.5\%$) and OOD performance by 1.1% (ICON, $93.8\%\rightarrow 94.9\%$).
AstroClassification.
@@
For AstroClassification, standard fine-tuning provides a significant performance boost over ERM: $71.6\%\rightarrow 78.9\%$ ID, $61.3\%\rightarrow 67.8\%$ OOD. ERM+targeted augs underperforms in ID accuracy compared to ERM alone ($71.6\%\rightarrow 68.8\%$) and standard fine-tuning ($78.9\%\rightarrow 68.8\%$), likely due to relatively strong targeted augmentations toward the redshift distribution of distant, faint objects present in the target distribution. However, OOD accuracy of ERM+targeted augs is competitive with standard fine-tuning, outperforming ERM. Self-training improves both ID and OOD performance compared to ERM but underperforms standard fine-tuning in both domains. Connect Later outperforms the best baseline, standard fine-tuning, by 12% OOD and 2% ID. The ID accuracy improves over standard fine-tuning despite the drop in ID accuracy from adding targeted augmentations to ERM, showing a complementary benefit between pretraining and targeted augmentations. Connect Later sets a new state-of-the-art OOD performance on AstroClassification by 3% over Avocado, a heavily tuned random forest model with expert-designed features [21].
Redshifts.
@@
Similarly to AstroClassification, standard fine-tuning significantly improves over ERM in both ID ($0.27\rightarrow 0.25$, 7% relative) and OOD ($0.32\rightarrow 0.28$, 13% relative) RMSE. Self-training performs similarly to ERM+targeted augs, while Connect Later outperforms the best baseline variant, standard fine-tuning, by 0.03 RMSE (11% relative) with comparable ID error. We use the same pretrained model for both AstroClassification and Redshifts for standard fine-tuning and Connect Later, demonstrating the reusability of pretrained representations.
@@ 5.2 Ablations
Number of Parameters ID Accuracy ($\uparrow$) OOD Accuracy ($\uparrow$)
21M (default) 78.47 79.49
69M 80.38 80.55
Connect Later ERM+targeted augs
ID Accuracy ($\uparrow$) OOD Accuracy ($\uparrow$) ID Accuracy ($\uparrow$) OOD Accuracy ($\uparrow$)
FT only 78.07 78.6 77.88 68.43
LP-FT 78.47 79.49 65.68 67.07
@@
We vary the strength of pretraining augmentations, which changes the connectivity between domains. This is most straightforward with the MAE objective, as augmentation strength is parameterized solely by masking percentage. We tested pretraining masking percentages  {20, 40, 60, 80, 90}while keeping the masking strategy unchanged (replace 10% of masked indices with random values from the lightcurve, another 10% are kept unchanged, and 80% are replaced with the mask token, which we choose to be 0). We show the ID and OOD test accuracy of each variant in Figure 3. Both ID and OOD performance peak at 60% masking, although we find that the performance of Connect Later is quite robust to the masking percentage, particularly for OOD performance. All of the masking percentages we tried improve on OOD performance over standard fine-tuning or ERM with targeted augmentations. Particularly, even with the strongest pretraining augmentations (90% masking), which should connect the domains more, the OOD performance did not improve over weaker augmentations. We hypothesize that increasing the strength of generic augmentations may indiscriminately increase the connectivity between all source and target examples, including examples from different classes that should not be strongly connected.
@@
Kumar et al. [28] showed that linear probing (with fixed neural embeddings) and then fine-tuning (LP-FT) the entire model improves both ID and OOD performance. Intuitively, full fine-tuning with a randomly initialized linear probe can destroy the pretrained features, and training the linear probe first mitigates this. We test LP-FT against FT only (all model weights are fine-tuned) with the Connect Later model and the ERM+targeted augs baseline. We find that LP-FT improves OOD accuracy by 0.9% over FT only when applied to Connect Later on AstroClassification (Table 6). On the other hand, LP-FT decreased OOD accuracy by 1.4% when applied to ERM+targeted augs, which uses random initialization (no pretraining). As a result, we use LP-FT on pretrained models but not on ERM or ERM+targeted augs.

### v2

@@ 4 Connect Later: Pretrain First, Targeted Augmentations Later
While our intuition about pretraining for domain adaptation stems from Shen et al. (2022), we show that applying targeted augmentations at fine-tuning time is sufficient for good generalization to the target domain even when the pretrained representations transfer across domains poorly. This allows us to reuse pretrained models for multiple downstream tasks.
@@ 4.1 Real-world examples of targeted augmentations
We design targeted augmentations for 4 real-world tasks: wildlife identification, tumor detection, and astronomical time-series classification and redshift prediction. We show examples from the source, augmented, and target datasets for these tasks in Figure 2.
Figure 2: Examples from the source dataset (left), an augmented version of the source example (middle), and the target dataset (right) for each of our tasks. (Top row) The AstroClassification and Redshifts tasks focus on time-varying astronomical objects observed in multiple wavelength ranges, plotted here as a multicolored time-series with each color corresponding to the wavelength range of the measurement. The redshifting augmentation simulates placing source objects at a higher redshift to better match the redshift distribution of the target dataset. The flux errors and flux values of the augmented example (middle) show much better resemblance to the target example. (Middle row) We randomize the habitat background by applying the Copy-Paste Same Y augmentation for iWildCam-WILDS (iWildCam-WILDS image examples shown here are from Gao et al. (2023)). (Bottom row) Stain Color Jitter alters the overall color of source images in Camelyon17-WILDS to improve performance on images from unseen hospitals.
@@
For iWildCam-WILDS (Beery et al., 2020; Sagawa et al., 2022), the task is to identify the wildlife species from static camera trap images. These cameras are placed in a wide variety of environments, which all have unique habitat conditions (e.g., African savannah vs. tropical rainforest) and camera characteristics (e.g., angles, resolutions).
@@
  • Targeted Augmentation: We augment the labeled dataset with the Copy-Paste Same Y algorithm, which uses image segmentation to copy-paste the animal onto different background images from cameras that have observed the same species (Gao et al., 2023).
@@
The task in Camelyon17-WILDS (Bandi et al., 2018) is to classify whether a patch of a histopathology slide contains a tumor. These slides are contributed from multiple hospitals, which use different stain colors and also vary in distributions of patient cancer stage.
@@
  • Targeted Augmentation: We augment the labeled dataset with the Stain Color Jitter algorithm, which jitters the color of the slide image in the hematoxylin and eosin staining color space (Tellez et al., 2018).
  • Task: Determine if a slide contains a tumor.
@@
Astronomical object classification (Boone, 2019; Allam Jr. & McEwen, 2022) involves predicting the object type (e.g., type II supernova) from a time series of an object’s brightness at multiple wavelengths. Expert labeling is only available for nearby objects, which are brighter and have different properties than distant objects (see Appendix A.1 for details).
@@
Similar to object type, redshift information is also available only for bright, nearby objects. We predict the scalar redshift value of each object and minimize mean squared error. Redshifts is a new dataset that we contribute as part of this work.
@@ 4.2 Designing targeted augmentations
Targeted augmentations offer the opportunity to incorporate domain knowledge to improve generalization performance. We provide a general methodology for designing targeted augmentations based on matching the target distribution on a feature space:
@@
    Identify a feature space $\mathcal{Z}$. We assume that we can label $z\in\mathcal{Z}$ for each input and that the source and target domains largely differ on this feature space. One such example is the space of spurious, domain-dependent features (e.g., camera angle or resolution for iWildCam-WILDS), which is the approach followed by Gao et al. (2023).
@@
We follow the procedure outlined above to design a targeted augmentation for AstroClassification and Redshifts (see Appendix B.2 for further details).
@@
    The source and target domains have different redshift distributions, so we identify this scalar feature as $z$.
@@
    We roughly fit the target redshift distribution within a reasonable range of the original redshift $z$, such that $\hat{p_{T}}(z^{\prime}\mid z)$ is distributed as $\text{loguniform}(0.95z,\;\text{min}(1.5(1+z)-1,\;5z))$, following Boone (2019).
@@ 5 Experiments
We empirically test Connect Later with contrastive pretraining (iWildCam-WILDS, Camelyon17-WILDS) as well as pretraining with masked autoencoding (AstroClassification, Redshifts) to demonstrate Connect Later as a general fine-tuning method. We note that masked autoencoding has been linked to contrastive learning in Zhang et al. (2022), which shows that the masked autoencoding objective upper bounds the contrastive loss between positive pairs – thus, masked autoencoding implicitly aligns the positive pairs induced by the masking augmentations.
@@
For iWildCam-WILDS, we use a ResNet-50 model pretrained on unlabeled ImageNet data with SwAV contrastive learning (Caron et al., 2020). We use a DenseNet121 pretrained on unlabeled data from Sagawa et al. (2022) with SwAV for Camelyon17-WILDS. We pretrain with masked autoencoding for AstroClassification and Redshifts by masking 60% of observations from each light curve (Appendix C). The same pretrained model is used for both tasks to demonstrate the reusability of pretrained features. We fine-tune the pretrained models with linear probing then fine-tuning (LP-FT, Kumar et al., 2022), which has been shown to improve OOD performance.
@@
We evaluate our framework against three baselines: ERM, ERM+targeted augs, and standard fine-tuning. We include Avocado (Boone, 2019), the state-of-the-art model for AstroClassification. We also include a self-training baseline for AstroClassification and Redshifts, which has been shown to perform well on some real-world datasets (Sagawa et al., 2022). For the self-training baseline, we pseudo-label the target dataset with a trained ERM+targeted augs model, then combine with the labeled source dataset and apply the targeted augmentation for training. We include additional domain adaptation baselines for iWildCam-WILDS and Camelyon17-WILDS: domain-adversarial neural networks (DANN, Ganin et al., 2016), correlation alignment (CORAL, Sun et al., 2016), Noisy Student (Xie et al., 2020b), and ICON [^1] https://github.com/a-tea-guy/ICON.
@@ 5.1 Main results
DANN  (Sagawa et al., 2022) $48.5\pm 3.2$ $31.9\pm 1.6$ $86.1\pm 1.3$ $64.5\pm 1.2$
CORAL  (Sagawa et al., 2022) $40.5\pm 1.6$ $27.9\pm 0.5$ $92.3\pm 0.7$ $62.3\pm 1.9$
Noisy Student  (Sagawa et al., 2022) $47.5\pm 1.0$ $32.1\pm 0.8$
@@
Tables 3 and 4 compare the results of Connect Later with baseline methods. Connect Later outperforms all baselines on the OOD metric while maintaining comparable or better ID performance and achieves state-of-the-art performance on iWildCam-WILDS by 0.8% OOD for ResNet-50, Camelyon17-WILDS by 1.1% OOD for DenseNet121, and AstroClassification by 3%.
Connect Later improves OOD performance when standard fine-tuning is minimally effective for UDA.
@@
On iWildCam-WILDS, standard fine-tuning minimally improves in OOD performance, while ERM+targeted augmentations improves by 6% ID and OOD over both ERM and standard fine-tuning. Connect Later improves over both standard fine-tuning (by $6.8\%$) and ERM+targeted augs (by $0.9\%$) in OOD performance, indicating that the pretrained representations as well as the targeted augmentations are both important for OOD performance. Connect Later achieves a new state-of-the-art performance for ResNet-50 on the iWildCam-WILDS benchmark.
When standard fine-tuning is effective, Connect Later still produces additional performance gains.
@@
On Camelyon17-WILDS, AstroClassification, and Redshifts, Connect Later still outperforms all variants even though standard fine-tuning already produces significant gains over ERM. For Camelyon17-WILDS, standard fine-tuning improves substantially over ERM in OOD average accuracy ($26.2\%$). ERM+targeted augs outperforms standard fine-tuning in ID accuracy by $4.4\%$, but does not improve OOD. Connect Later sets a new state-of-the-art on Camelyon17-WILDS with DenseNet121, improving on the best ID performance by 1.8% (ERM+targeted augs) and OOD performance by 1.1% (ICON).
@@
For AstroClassification, standard fine-tuning also performs significantly better than ERM: $7\%$ ID, $6.5\%$ OOD. ERM+targeted augs underperforms in ID accuracy compared to ERM ($-2.8\%$) and standard fine-tuning ($-9.9\%$), likely due to the strong targeted augmentations. However, OOD accuracy of ERM+targeted augs is competitive with standard fine-tuning, outperforming ERM. Connect Later outperforms the best baseline, standard fine-tuning, by 12% OOD and 2% ID. The ID accuracy outperforms both standard fine-tuning and ERM+targeted augs, showing a complementary benefit between pretraining and targeted augmentations. Connect Later sets a new state-of-the-art OOD performance on AstroClassification by 3% over Avocado, a heavily tuned random forest model with expert-designed features (Boone, 2019).
@@
Redshifts results are similar to AstroClassification, with standard fine-tuning significantly improving over ERM in both ID (7% relative) and OOD (13% relative) RMSE. Connect Later outperforms the best baseline variant, standard fine-tuning, by 0.03 RMSE (11% relative) with comparable ID error.
Connect Later improves OOD performance for CLIP fine-tuning.
We additionally evaluated the effectiveness of Connect Later for CLIP ViT-L/14 (Radford et al., 2021) on the iWildCam-WILDS dataset (Table 5). Standard fine-tuning improves substantially over both ERM and ERM + targeted augs, likely due to CLIP’s internet-scale pretraining dataset as well as the practical importance of pretraining for ViT performance. However, Connect Later still delivers additional gains over standard fine-tuning (0.5% ID, 1.4% OOD).
Table 5: ID and OOD Macro F1 results for fine-tuning CLIP ViT-L on iWildCam-WILDS. Results are averaged over 15 trials and we report 95% confidence intervals on each mean estimate. Rows with means within 1 interval of the best mean are bolded.
ID Test OOD Test
ERM $22.6\pm 0.6$ $7.8\pm 0.2$
Standard fine-tuning $\mathbf{55.3\pm 1.4}$ $42.8\pm 0.9$
ERM + targeted augs $23.8\pm 0.7$ $8.4\pm 0.4$
Connect Later $\mathbf{55.8\pm 0.8}$ $\mathbf{44.2\pm 0.9}$
Other baselines.
DANN, CORAL, and Noisy Student did not produce competitive OOD average accuracy for either iWildCam-WILDS or Camelyon17-WILDS. ICON is the best baseline for Camelyon17-WILDS OOD average accuracy and is outperformed only by Connect Later. For AstroClassification and Redshifts, self-training improves both ID and OOD performance compared to ERM but underperforms standard fine-tuning in both domains.
@@ 5.2 Ablations
We vary the strength of pretraining augmentations with the MAE objective, as augmentation strength is parameterized solely by masking percentage. We tested pretraining masking percentages  {20, 40, 60, 80, 90}with the same masking strategy (replace 10% of masked indices with random values from the time-series, another 10% are kept unchanged, and 80% are replaced with the mask token, 0). We show the ID and OOD test accuracy of each variant in Figure 3. Both ID and OOD performance peak at 60% masking, although the performance of Connect Later is quite robust to the masking percentage. All masking percentages improve on OOD performance over standard fine-tuning or ERM with targeted augmentations. Even the strongest augmentations (90% masking) did not improve OOD performance over weaker augmentations. We hypothesize that strong generic augmentations may indiscriminately increase the connectivity between all source and target examples, including examples from different classes that should not be strongly connected.
Number of Parameters ID Acc ($\uparrow$) OOD Acc ($\uparrow$)
21M (default) 78.47 79.49
69M 80.38 80.55
Connect Later ERM+targeted augs
ID Acc($\uparrow$) OOD Acc($\uparrow$) ID Acc($\uparrow$) OOD Acc($\uparrow$)
FT only 78.07 78.6 77.88 68.43
LP-FT 78.47 79.49 65.68 67.07
@@
Kumar et al. (2022) showed that linear probing (with fixed neural embeddings) and then fine-tuning (LP-FT) the entire model improves both ID and OOD performance. Intuitively, full fine-tuning with a randomly initialized linear probe can destroy the pretrained features, and training the linear probe first mitigates this. We test LP-FT against FT only (all model weights are fine-tuned) with the Connect Later model and the ERM+targeted augs baseline. We find that LP-FT improves OOD accuracy by 0.9% over FT only when applied to Connect Later on AstroClassification (Table 7). On the other hand, LP-FT decreased OOD accuracy by 1.4% when applied to ERM+targeted augs, which uses random initialization (no pretraining). As a result, we use LP-FT on pretrained models but not on ERM or ERM+targeted augs.